OpenStudy (anonymous):

Shades R Us charges $20 per day to rent a lounge chair and $15 per day to rent an umbrella. Dan and Lisa paid a total of $245 to rent a lounge chair and an umbrella each during their vacation. Lisa rented the chair and umbrella for 1 day less than Dan. The following equation represents this situation, where x represents the number of days Dan rented the lounge chair and umbrella. 20x + 15x + 20(x − 1) + 15(x − 1) = 245 What is the total amount Dan paid to rent the lounge chair and umbrella during his vacation?

OpenStudy (texaschic101):

20x + 15x + 20(x - 1) + 15(x - 1) = 245 35x + 20x - 20 + 15x - 15 = 245 35x + 20x + 15x = 245 + 20 + 15 70x = 280 x = 280/70 x = 4 Dan rented for 4 days... 20*4 = 80 15 * 4 = 60 total = 140 Lisa rented for 3 days 20 * 3 = 60 15 * 3 = 45 total = 105 check.... Dan spent (20 * 4) + (15 * 4) = 140 <-- your answer Lisa spent (20 * 3) + (15 * 3) = 105 totaling (140 + 105) = 245